# Artifacts or Abduction: How Do LLMs Answer Multiple-Choice Questions Without the Question?

### v1

@@ Abstract
Multiple-choice question answering (MCQA) is often used to evaluate large language models (LLMs). To see if MCQA assesses LLMs as intended, we probe if LLMs can perform MCQA with choices-only prompts, where models must select the correct answer only from the choices. In three MCQA datasets and four LLMs, this prompt bests a majority baseline in 11/12 cases, with up to 0.33 accuracy gain. To help explain this behavior, we conduct an in-depth, black-box analysis on memorization, choice dynamics, and question inference. Our key findings are threefold. First, we find no evidence that the choices-only accuracy stems from memorization alone. Second, priors over individual choices do not fully explain choices-only accuracy, hinting that LLMs use the group dynamics of choices. Third, LLMs have some ability to infer a relevant question from choices, and surprisingly can sometimes even match the original question. We hope to motivate the use of stronger baselines in MCQA benchmarks, the design of robust MCQA datasets, and further efforts to explain LLM decision-making. [^1] Our data and code is available at https://github.com/nbalepur/mcqa-artifacts
@@ 1 Introduction
AQI performs similarly to the choices-only prompt, even besting it for 3 LLMs on HellaSwag. The two strategies exhibit moderate agreement, implying that AQI may be just one of many strategies LLMs use in choices-only prompts. We then study the questions generated by LLaMA on ARC. Out of the 43 cases when AQI picks the gold answer and the inferred question is answerable by one of the choices, 42% match the meaning of the original question. Thus, we argue that a successful choices-only prompt does not always mean that the LLM uses surface-level shortcuts linked with artifacts; in some cases, it may recover the missing question.
@@ 2.3 Datasets
We select three MCQA benchmarks from the Open LLM Leaderboard Beeching et al. (2023): ARC-Challenge Clark et al. (2018), MMLU Hendrycks et al. (2021), and HellaSwag Zellers et al. (2019), collectively testing factual knowledge, scientific reasoning, and commonsense reasoning. We follow the Open LLM Leaderboard to pick few-shot examples and use 25 random training examples for ARC, 10 random training examples for HellaSwag, and the 5 given training examples for each subject in MMLU. We use the entire evaluation sets, but omit the questions in ARC that do not have four choices to ensure all questions have four choices.
@@ 2.4 Evaluation Protocol
Figure 3: LLM accuracy with full prompts versus our three tested memorization prompts
@@ 3 Performing MCQA With No Question
To test if LLMs exploit MCQA artifacts, we design partial-input choices-only prompts (section˜3.1), and show that these prompts often largely best majority baselines (section˜3.2). High partial-input accuracies typically imply artifact exploitation via simple shortcuts Du et al. (2023). However, in §6 we explore if some of this overperformance can be attributed to the LLM’s ability to infer the original question. In this case, it would be difficult to claim that the model has bypassed the intended MCQA format. Thus, in this section, we establish an initial ceiling for artifact exploitation in the choices-only setting.
@@ 3.2 Results
In 11/12 cases, the choices-only prompt surpasses the majority baseline significantly, indicating that LLMs may be using artifacts in MCQA benchmarks (Figure 2). Further, larger LLMs tend to have higher choices-only accuracy, implying that more capable LLMs may use artifacts to a higher degree. In Appendix B.1, we analyze scaling laws with LLaMA to understand this relation more.
@@ 6.1 Prompt Design
We extract $q_{gen}$ from Prompt 6.1444The LLM also outputs $\texttt{Answer:}\;a$” after $q_{gen}$. We first wanted AQI to be 1 step where $q_{gen}$ acts as a chain-of-thought, but this 1-step process is less effective (see Appendix B.5) to perform step two of AQI, where the LLM answers $q_{gen}$:
@@ 6.2 Quantitative Results
AQI always results in an accuracy near the choices-only prompt on ARC and MMLU (Figure 5). One hypothesis is that the LLMs ignore the generated question $q_{gen}$ in the self-ask prompt (Prompt 6.1), but in 11/12 cases, using the random question $q_{rand}$ underperforms $q_{gen}$. Thus, $q_{gen}$ is of higher average quality than $q_{rand}$, meaning our LLMs can extract information from $q_{gen}$ relevant to answering the question. Further, on HellaSwag for Falcon, Phi, and Mixtral, AQI bests the choices-only prompt. AQI is a reasoning-based strategy and thus, future works can try to reason with other strategies Huang and Chang (2023) to see if LLMs can obtain even higher accuracy in choices-only settings.
@@
Table 1: Cohen’s $\kappa$ for the correctness of the choices-only prompt (Prompt 3.1) and AQI (Prompts 6.1, 6.1). Random is the average Cohen’s $\kappa$ of the choices-only prompt and a random uniform distribution.
@@ 8 Conclusion
We find that LLMs can perform MCQA without access to the question, an ability elicited by minimal few-shot examples. In pursuing explanations for these results, we release an in-depth black-box evaluation suite, based on our proposed partial-input prompts. As part of our analysis, we show LLMs still exhibit reasoning abilities in choices-only settings via the meta-strategy of question inference. Along with software and analysis, we suggest ways to evaluate LLMs on MCQA as intended, where both the question and choices are needed, including: 1) reporting stronger baselines like our proposed choices-only prompt; and 2) designing robust data creation protocols. In light of our findings, we advocate for more critical discussions around LLM evaluation, allowing us to understand the capabilities our benchmarks are truly assessing.
@@ 9 Limitations
Lastly, LLMs are sensitive to prompts Min et al. (2022) and hyperparameters, so it is possible that varied configurations could alter the accuracy in our experiments. Our prompts and hyperparameters were just the first ones we tried, so we believe that LLMs can likely be optimized to exploit MCQA artifacts to an even higher degree. Future works can explore these alterations more in-depth and analyze LLM sensitivity in choices-only settings.
@@ 10 Ethical Considerations
Through few-shot choices-only prompts, we discover that LLMs can achieve high choices-only accuracy in prominent MCQA benchmark datasets. However, our goal is not to discount the adoption of these datasets for evaluation and suggest that they are fundamentally flawed, or discredit the performance of LLMs on these datasets and claim that they are incapable of performing MCQA. Instead, we wish to provide more transparency for MCQA benchmarking and LLM decision-making, which we hope will inspire future work on: 1) better understanding LLM decision-making in MCQA; 2) reporting choices-only baselines as a stronger alternative to majority class baselines on MCQA evaluation suites; and 3) designing more resilient benchmarks that limit the influence of artifacts.
@@ B.6.2 Additional Results
Finally, we provide examples of questions generated from AQI. In Tables 6 and 7, we show generated questions when inferring the question succeeds, and the generated question matches and does not match the original question, respectively. In Table 8, we show generated questions when inferring the question fails, but the question is still answerable.

### v2

@@ Abstract
Multiple-choice question answering (MCQA) is often used to evaluate large language models (LLMs). To see if MCQA assesses LLMs as intended, we probe if LLMs can perform MCQA with choices-only prompts, where models must select the correct answer only from the choices. In three MCQA datasets and four LLMs, this prompt bests a majority baseline in 11/12 cases, with up to 0.33 accuracy gain. To help explain this behavior, we conduct an in-depth, black-box analysis on memorization, choice dynamics, and question inference. Our key findings are threefold. First, we find no evidence that the choices-only accuracy stems from memorization alone. Second, priors over individual choices do not fully explain choices-only accuracy, hinting that LLMs use the group dynamics of choices. Third, LLMs have some ability to infer a relevant question from choices, and surprisingly can sometimes even match the original question. Inferring the original question is an impressive reasoning strategy, but it cannot fully explain the high choices-only accuracy of LLMs in MCQA. Thus, while LLMs are not fully incapable of reasoning in MCQA, we still advocate for the use of stronger baselines in MCQA benchmarks, the design of robust MCQA datasets for fair evaluations, and further efforts to explain LLM decision-making. [^1] Our data and code are available at https://github.com/nbalepur/mcqa-artifacts
@@ 1 Introduction
AQI performs similarly to the choices-only prompt, even besting it for 3 LLMs on HellaSwag, and they also exhibit moderate agreement, implying that AQI may be just one of many strategies LLMs use in choices-only settings. We then study the questions generated by LLaMA on ARC. Out of the 43 cases when AQI picks the gold answer and the inferred question is answerable by one of the choices, 42% match the meaning of the original question. Thus, we argue that a successful choices-only prompt does not always mean that the LLM uses surface-level shortcuts linked with artifacts; in some cases, it may recover the missing question.
@@ 2.3 Datasets
We select three MCQA benchmarks from the Open LLM Leaderboard Beeching et al. (2023): ARC-Challenge Clark et al. (2018), MMLU Hendrycks et al. (2021), and HellaSwag Zellers et al. (2019), collectively testing factual knowledge, scientific reasoning, and commonsense reasoning. We use the entire evaluation sets, but omit the questions in ARC that do not have four choices to ensure all questions have four choices. We follow the Open LLM Leaderboard to pick few-shot examples and use 25 random training examples for ARC, 10 random training examples for HellaSwag, and the 5 given training examples for each subject in MMLU. While our experiments reflect the prompting setups of present MCQA benchmarks, we believe that future works could extend our analysis with zero-shot prompting Kojima et al. (2022) to limit the influence of priors from the few-shot examples.
@@ 2.4 Evaluation Protocol
Figure 3: LLM accuracy with full prompts versus our three tested memorization prompts. An asterisk (*) denotes that the choices-only prompt significantly outperforms the majority class baseline (two-sample $t$-test, $p<5\text{e-}5$).
@@ 3 Performing MCQA With No Question
To test if LLMs exploit MCQA artifacts, we design partial-input choices-only prompts (section˜3.1), and show that these prompts often largely best majority baselines [^4] A majority class baseline always predicts the most frequent answer choice found in the dataset. (section˜3.2). High partial-input accuracies typically imply artifact exploitation via simple shortcuts Du et al. (2023). However, in §6 we explore if some of this overperformance can be attributed to the LLM’s ability to infer the original question. In this case, it would be difficult to claim that the model has bypassed the intended MCQA format. Thus, in this section, we establish an initial ceiling for artifact exploitation in the choices-only setting.
@@ 3.2 Results
In 11/12 cases, the choices-only prompt surpasses the majority baseline significantly, indicating that LLMs may be using artifacts in MCQA benchmarks (Figure 2). Further, larger LLMs tend to have higher choices-only accuracy, implying that more capable LLMs may use artifacts to a higher degree. In Appendix B.1, we study scaling laws with LLaMA to understand this relation more. Overall, these results serve as an initial ceiling for artifact exploitation in choices-only settings, which we use to motivate our subsequent analyses that help attribute where this accuracy may stem from.
@@ 6.1 Prompt Design
We extract $q_{gen}$ from Prompt 6.1555The LLM also outputs $\texttt{Answer:}\;a$” after $q_{gen}$. We initially wanted AQI to be 1 step where $q_{gen}$ acts as a chain-of-thought (i.e. merge steps one and two of AQI), but this 1-step process is less effective (see Appendix B.5). to perform step two of AQI, where the LLM answers $q_{gen}$:
@@ 6.2 Quantitative Results
AQI always results in an accuracy near the choices-only prompt on ARC and MMLU (Figure 5). One hypothesis is that the LLMs ignore the generated question $q_{gen}$ in the self-ask prompt (Prompt 6.1), but in 11/12 cases, using the random question $q_{rand}$ underperforms $q_{gen}$. Thus, $q_{gen}$ is of higher average quality [^6] It is also possible for $q_{rand}$ to be misleading and $q_{gen}$ to just be irrelevant/non-misleading. However, our qualitative analysis (section 6.3) reveals that $q_{gen}$ is often answerable by the choices, and thus we claim that it is higher quality than $q_{rand}$. than $q_{rand}$, meaning our LLMs can extract information from $q_{gen}$ relevant to answering the question. Further, on HellaSwag for Falcon, Phi, and Mixtral, AQI bests the choices-only prompt. AQI is a reasoning-based strategy and thus, future works can try to reason with other strategies Huang and Chang (2023) to see if LLMs can obtain even higher accuracy in choices-only settings.
@@
Table 1: Cohen’s $\kappa$ for the correctness of the choices-only prompt (Prompt 3.1) and AQI (Prompts 6.1, 6.1). Random is the average Cohen’s $\kappa$ of the choices-only prompt and a random uniform distribution. Appendix Table 5 shows the same agreement for 1-step AQI.
@@ 8 Conclusion
We find that LLMs can perform MCQA without access to the question, a result that can be achieved through few-shot prompting with minimal examples. In pursuing explanations for these results, we release an in-depth black-box evaluation suite, based on our proposed partial-input prompts. As part of our analysis, we show that LLMs still exhibit reasoning abilities in choices-only settings via the meta-strategy of question inference, suggesting that partial-input abilities of LLMs should not be completely attributed to surface-level shortcuts linked to artifacts. Along with software and analysis, we suggest ways to evaluate LLMs on MCQA as intended, where both the question and choices are needed, including: 1) reporting stronger baselines like our proposed choices-only prompt; and 2) designing robust data creation protocols. In light of our findings, we advocate for more critical discussions around LLM evaluation, allowing us to better understand and interpret the model capabilities our benchmarks are truly assessing.
@@ 9 Limitations
Lastly, LLMs are sensitive to prompts Min et al. (2022) and hyperparameters, so it is possible that varied configurations could alter the accuracy in our experiments. Our prompts and hyperparameters were just the first ones we tried, so we believe that LLMs can likely be optimized to obtain even higher choices-only accuracy. Future works can explore these alterations more in-depth and analyze LLM sensitivity in choices-only settings.
@@ 10 Ethical Considerations
Through few-shot prompting, we discover that LLMs can achieve high choices-only accuracy in prominent MCQA benchmark datasets. However, our goal is not to discount the adoption of these datasets for evaluation and suggest that they are fundamentally flawed, or discredit the performance of LLMs on these datasets and claim that they are incapable of performing MCQA. Instead, we wish to provide more transparency for MCQA benchmarking and LLM decision-making, which we hope will inspire future work on: 1) better understanding LLM decision-making in MCQA; 2) reporting choices-only baselines as a stronger alternative to majority class baselines on MCQA evaluation suites; and 3) designing more resilient benchmarks that limit the influence of artifacts.
11 Acknowledgements
We would like to thank members of the CLIP lab at the University of Maryland and external collaborators for their feedback and discussions of this work, including Yu (Hope) Hou, Dayeon (Zoey) Ki, Neha Srikanth, Rupak Sarkar, Shi Feng, and Jordan Boyd-Graber. We are also grateful for our discussions with Hailey Schoelkopf, Stella Biderman, and other members of the EleutherAI community. We also thank the anonymous reviewers for their feedback. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2236417. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ B.6.2 Additional Results
Finally, we provide examples of questions generated from AQI. In Tables 7 and 8, we show generated questions when inferring the question succeeds, and the generated question matches and does not match the original question, respectively. In Table 9, we show generated questions when inferring the question fails, but the question is still answerable.
@@ Appendix C Related Work on MCQA Evaluations
Dataset LLaMA Falcon Mistral Phi
ARC 0.487 0.628 0.495 0.459
MMLU 0.194 0.221 0.293 0.129
HellaSwag 0.415 0.563 0.557 0.458
Table 5: Cohen’s $\kappa$ for the correctness of the choices-only prompt (Prompt 3.1) and 1-step AQI. The average $\kappa$ is 0.41